OpenStudy (anonymous):

how do you prove?? sinx/1+cosx=1-cosx/sinx

hartnn (hartnn):

@rl88 , still need help with this ?

hartnn (hartnn):

First, we multiply numerator and denominator by 1-cos x \(\large \dfrac{\sin x}{1+\cos x}=\dfrac{\sin x}{1+\cos x}\dfrac{1-\cos x}{1-\cos x}\) now denominator is of the form \((a+b)(a-b)\) which simplifies to \(a^2-b^2\) so, \((1-\cos x)(1+\cos x)=...?\)